• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Ole Hansen

    Log entry time 16:06:14 on April29,2008

    Entry number 235801

    keyword=Day shift summary

    Kinematics E
    4/29/08

    run events mC sum mC comment

    4369 1M, 2.686mC 75.098 Al 6cm dummy
    4370 1M 1.351 76.449
    4371 1M 1.368 77.817
    4372 1M 1.346 79.163
    4373 1M 1.347 80.51
    4374 1M 1.352 81.862
    4375 1M 1.372 83.234
    4376 0.9 84.134
    4377 1M 1.353 85.487 BB HV trip
    4378 1M 1.356 86.843
    4379 1M 1.362 88.205
    4380 0.321 88.526 end for beam studies

    Beam studies from 10am. Original ETA 2pm. Still no beam at end of shift.

    Took opportunistic controlled access from about 10:30am - 3:30am to move spectrometers, refill BigBite He bags, and diagnose Moller problems (sucessfully).

    Runs 4381-4386 are DAQ tests (see halog), taken during beam studies.